# Rank, Chunk, and Expand: Lineage-Oriented Reasoning for Taxonomy Expansion

### v1

@@ 5.4 Case Studies
Table 3 discusses case studies for "Env" and "Sci" datasets, highlighting the reasoning mechanisms of LORex’s components. We observe that parent retriever mostly reasons over paths to directly select the most appropriate parent from candidate terms, as in the case of “Arctic Ocean,” “pteridology,” and “realpolitik.” The path-based verifier effectively exploits paths and reasons over them to distinguish between the retrieved path and alternatives, which is clearly evident in the case of “pteridology,” where the best-inferred path is “pteridology > botany > biology > life science > natural science > science.” In contrast, the candidate-based verifier relies only on definitions to verify the retrieved candidate, as seen in the case of “realpolitik,” where it fails to capture the appropriate hypernym. These results emphasize the importance of path-based reasoning in verification over candidate-based verification, which introduces unnecessary abstraction. More such cases are discussed in Appendix K.
@@ 6 Conclusion
While LORex effectively retrieves parent terms in taxonomy expansion, its performance heavily relies on the quality of the candidate ranking. If the correct parent term is not highly ranked, LORex suffers from increased retrieval errors and higher computational overhead due to multiple iterations of retrieval and verification. This iterative process significantly impacts both latency and accuracy, particularly in cases where the ranker fails to surface relevant candidates early. As a result, the overall system efficiency is constrained by the discriminative ranker, making LORex highly sensitive to ranking quality. In other words, “the method is only as good as its ranker.” Exploring better ranking techniques and reducing the reliance of retrievers on rankers will be the focus of our future works. We will also focus on improving interpretability by employing an ensemble of LLMs to reason over hierarchies.
@@
Appendix A Euler Tour
As discussed in Section 4.1, we present an example of an Euler-toured path for taxonomy shown in Figure 6. The Euler toured path is defined as $P$ = ["water pollution", "pollution", "environment", "pollution", "air pollution", "pollution", "soil pollution", "pollution", "water pollution", "marine pollution", "water pollution", "chemical pollution"]. Therefore, the two types of verbalized paths (with and without query definition are discussed in Box A where $P\in P_{v}(q,P)$ while $P_{r}\in P_{v}(P)$. These paths are used to compute three types of losses as shown in Figure 7.
@@ Appendix A Euler Tour
Box A: Verbalized Paths $P$ potable water pollution is the contamination of water that is intended for human consumption [SEP] "water pollution" is child of "pollution" is child of "environment" is parent of "pollution" is parent of "air pollution" is child of "pollution" is parent of "soil pollution" is child of "pollution" is parent of "water pollution" is parent of "marine pollution" is child of "water pollution" is parent of "chemical pollution" $P_{r}$water pollution" is child of "pollution" is child of "environment" is parent of "pollution" is parent of "air pollution" is child of "pollution" is parent of "soil pollution" is child of "pollution" is parent of "water pollution" is parent of "marine pollution" is child of "water pollution" is parent of "chemical pollution"
@@ Appendix B Semantic Filter Prompt
The semantic filter module, discussed in Section 4.2, relies on LLM to check if the chunk of candidates is relevant. The prompt example is discussed in Prompt  B.1.
Prompt B.1: Semantic Filter Prompt Example You are a semantic relevance expert for terms present in a taxonomy. Your task is to determine whether the set of candidate terms is a semantically relevant match for the given query term ’Arctic Ocean’ and ’Environment’ taxonomy. List of Candidate terms: - geophysical environment - ocean - wild mammal - animal life - climatic zone Definitions of candidate terms: - geophysical environment - environment most often refers to: - ocean - an ocean is a body of water that composes much of a planet’s hydrosphere - wild mammal - mammals, fur or hair, and three middle ear bones - animal life - animals are multicellular eukaryotic organisms that form the biological kingdom animalia - climatic zone - climate classification systems are ways of classifying the world’s climates Reason over the definitions of candidate terms to determine their relevance as a whole with respect to query terms Arctic Ocean and Environment taxonomy. If at least one of the candidate terms is the most relevant parent term for the taxonomy, return YES. Otherwise, return NO. Answer:
@@ Appendix C Parent Retrieval Prompt
The parent retriever module, discussed in Section 4.3, reasons over taxonomy paths to determine the appropriate parent from the chunk of candidate terms. An example of the parent retriever prompt is discussed in Prompt  C.1.
Prompt C.1: Parent Retriever Prompt Example You are an expert in hypernymy (is-a) relationship detection for a taxonomy. Your task is to find the most appropriate candidate hypernym of the query node ’Arctic Ocean’ within the ’environment’ taxonomy. The most appropriate hypernym is the most granular category that directly encompasses the query term. List of Candidate terms: - geophysical environment - ocean - wild mammal - animal life - climatic zone Definitions of candidate terms: - geophysical environment - environment most often refers to: - ocean - an ocean is a body of water that composes much of a planet’s hydrosphere - wild mammal - mammals, fur or hair, and three middle ear bones - animal life - animals are multicellular eukaryotic organisms that form the biological kingdom animalia - climatic zone - climate classification systems are ways of classifying the world’s climates Path from query term to root node for all candidate terms: - Arctic Ocean -> geophysical environment -> environment - Arctic Ocean -> ocean -> geophysical environment -> environment - Arctic Ocean -> wild mammal -> animal life -> wildlife -> environment - Arctic Ocean -> animal life -> wildlife -> environment - Arctic Ocean -> climatic zone -> climate -> environment Query Node Definition: The Arctic Ocean is the smallest and shallowest of the world’s oceans. Some examples of hypernymy relationships in the taxonomy are as follows: - Children of geophysical environment are: arid zone, estuary, island, lake, mountain, ocean, plain, polar region, sea - Parent of animal life is: wildlife - Children of climatic zone are: equatorial zone, frigid zone, humid zone, subtropical zone, temperate zone, tropical zone - Parent of climatic zone is: climate Instructions: - Determine the most appropriate direct hypernym of ’Arctic Ocean’ from the given list of candidate terms. Do not return any term which is not in the list. - A hypernym must be a most granular category in which the query node is an instance. - If no candidate term correctly fits as a hypernym , return NOT FOUND. - Do not include explanations, justifications, or additional context. Answer:
@@ Appendix D Parent Verifier Prompt
The parent verifier module verifies the appropriateness of the retrieved candidate by reasoning on paths as discussed in Section 4.4. An example of the prompt is discussed in Prompt  D.1.
Prompt D.1: Parent Verifier Prompt Example You are an expert verifier of hypernymy relationship for a taxonomy using paths. You have been given the following path from query term ’Arctic Ocean’ to root node: Arctic Ocean -> ocean -> geophysical environment -> environment Your task is to verify if the given path is the most appropriate path for the query node or the following paths provide a better alternative. Other possible paths: - Arctic Ocean -> geophysical environment -> environment - Arctic Ocean -> ocean -> geophysical environment -> environment - Arctic Ocean -> wild mammal -> animal life -> wildlife -> environment - Arctic Ocean -> animal life -> wildlife -> environment - Arctic Ocean -> climatic zone -> climate -> environment Some examples of hypernymy relationships in the taxonomy are as follows: - Children of geophysical environment are: arid zone, estuary, island, lake, mountain, ocean, plain, polar region, sea - Parent of animal life is: wildlife - Children of climatic zone are: equatorial zone, frigid zone, humid zone, subtropical zone, temperate zone, tropical zone - Parent of climatic zone is: climate Definitions of candidate terms: - geophysical environment - environment most often refers to: - ocean - an ocean is a body of water that composes much of a planet’s hydrosphere - wild mammal - mammals, fur or hair, and three middle ear bones - animal life - animals are multicellular eukaryotic organisms that form the biological kingdom animalia - climatic zone - climate classification systems are ways of classifying the world’s climates Query Node Definition: The Arctic Ocean is the smallest and shallowest of the world’s oceans. Instructions: - Compare the given path with the alternative paths based on definitions and hypernymy relationships. - Select the most appropriate path that best represents the hierarchical relationship of ’Arctic Ocean’ in the taxonomy. Return the most appropriate path. Answer:
Appendix E Pseudocode
@@ 14      $\mathcal{B}\leftarrow\mathcal{B}\setminus B^{i}_{j}$;
Hyperparameter Sci Food Env WordNet
Batch Size 32 32 64 32
Learning Rate 2e-5 1e-5 2e-5 1e-5
Max Padding 350 350 200 350
$d$ 0.2 0.2 0.2 0.2
$\epsilon$ 1e-8 1e-8 1e-8 1e-8
Epochs 20 50 20 20
Table 5: Hyperparameter settings for different datasets.
@@ Appendix F Implementation Details
LORex  and BERT+MLP are implemented in PyTorch, with all other baselines obtained from their respective official repositories. The training and inference processes are executed on a single 80GB NVIDIA A100 GPU to ensure computational efficiency. To implement TEMPORA, we use bert-base-uncased as the default pre-trained model, optimized using AdamW. The hyperparameters utilized for training and inference are detailed in Table 5. Notably, TEMPORA is trained on all benchmarks for fewer epochs compared to TEMP, as the primary objective is to achieve a ranker that delivers sufficiently high performance without excessive fine-tuning. Retrieval and verification rely on instruction-tuned models from the LLaMA family, namely LLaMA-3.1-8B-Instruct, LLaMA-3-8B-Instruct, and LLaMA-2-7B-Chat. As demonstrated in Figure 8, all benchmarks except Food achieve a Hit@15 score above 0.9, indicating that for 90% of queries the correct parent exists within the top-15 candidates. Therefore, to enhance the efficiency of retrieval and verification, inference is performed on three chunks or 15 candidates simultaneously. If any chunk is empty, it is discarded, and only the remaining chunks are considered. This strategy ensures that the correct output is obtained in a single inference pass, optimizing both computational efficiency and retrieval accuracy. For text generation tasks, a stopping criterion is enforced to terminate generation upon encountering a newline character, which prevents the generation of unnecessary tokens, thereby reducing inference latency.
@@ Appendix H Baseline Methods
  • FLAME Mishra et al. (2024) fine-tunes low-rank parameters of LLMs to predict parent terms by leveraging semantically relevant local and global neighborhood information of candidate nodes.
Appendix I Evaluation Metrics
    where $\mathbb{I}(\cdot)$ denotes the indicator function. Wu & Palmer Similarity (Wu&P): This metric quantifies the structural similarity between the predicted parent and the ground truth within the seed taxonomy. It is computed based on the closest common ancestor,
[…]
Method Food WordNet
Hit@1 Hit@10 Hit@1 Hit@10
BM25 0.071 0.165 0.065 0.171
SBERTCOS 0.182 0.380 0.113 0.390
BERTCOS 0.077 0.118 0.076 0.225
SimCSE 0.119 0.280 0.258 0.621
TEMP 0.330 0.529 0.329 0.682
TEMPORA 0.451 0.663 0.405 0.794
Appendix J Discriminator Results
Appendix K Case Studies
Extending on the case studies discussed in Section 5.4, we further analyze some additional examples with more semantically rich explanations in Table 8. The parent retriever provides straightforward reasoning over definitions and path-based relationships. However, as it is susceptible to errors, we rely on the verifier’s reasoning to validate and refine the retriever’s predictions. The Path-Based Verifier employs connectivity-based reasoning to assess retrieval accuracy by analyzing hierarchical paths. For instance, in the case of ‘grapefruit peel’, the verifier explains, “The given path is the most appropriate as it directly connects ‘grapefruit peel’ to the root node ‘food’ through a series of hypernymy relationships that align with its definition.” In contrast, the candidate-based verifier relies on definitions, which can lead to incorrect predictions when contextual hierarchical relationships are not explicitly captured in definitions.
Query Parent Retriever Path-Based Verifier Candidate-Based Verifier
@@ Appendix I Evaluation Metrics
candied citrus peel
@@
nominal aphasia
Inability to name objects or recognize written or spoken names of objects.
aphasia
The most appropriate direct hypernym of ’nominal aphasia’ is ’aphasia.’
nominal aphasia > aphasia > brain disorder
The given path is the most appropriate path for the query node ’nominal aphasia’. This is because ’nominal aphasia’ is a type of ’aphasia,’ which is a type of ’brain disorder.’ The path ’nominal aphasia -> aphasia -> brain disorder’ accurately represents the hierarchical relationship of ’nominal aphasia’ in the taxonomy. The other alternative paths do not accurately represent the hierarchical relationship of ’nominal aphasia’ in the taxonomy.
agnosia
The query node ’nominal aphasia’ is a type of aphasia which is a brain disorder. However, the query node is more specifically related to the inability to recognize objects by use of the senses which is the definition of agnosia. Therefore, agnosia is a more specific and appropriate parent for the query node ’nominal aphasia’. The given candidate ’aphasia’ is a broader term that encompasses various types of aphasia including nominal aphasia. However, it does not specifically capture the essence of the query node.
@@
Table 8: Case study analyzing the impact of LORex’s components on "Food" and "WordNet" benchmarks. The table presents component outputs in bold, with their definitions and explanations in italics. Yellow highlighted text denotes key phrases in definitions and explanations that play an important role in reasoning.

### v4

@@ 5.4 Case Studies
Table 6 discusses case studies for "Env" and "Sci" datasets, highlighting the reasoning mechanisms of LORex’s components. We select two samples of each dataset and observe that parent retriever mostly reasons over paths to directly select the most appropriate parent from candidate terms as visible in case of query terms “Arctic Ocean” and “exhaustion of resources.” However, it fails to retrieve the correct hypernym in case of “pteridology” whose definition does not clearly define the term, creating confusion for parent retriever. Moreover, the path-based verifier effectively exploits paths and reasons over them to distinguish between the retrieved path and alternatives, which is clearly evident in the case of “pteridology,” where the best-inferred path is “pteridology > botany > biology > life science > natural science > science.” In contrast, the candidate-based verifier relies only on definitions, which are noisy, incomplete and incorrect, to verify the retrieved candidate, as seen in the case of “realpolitik,” where it fails to capture the appropriate hypernym, giving “social science” as the hypernym. These results emphasize the importance of path-based reasoning in verification after candidate retrieval over verification using definitions, which introduces unnecessary abstraction. More such cases are discussed in Appendix L.
@@ 6 Conclusion
Acknowledgements
We gratefully acknowledge the support of Microsoft through the Microsoft Academic Partnership Grant 2024 (MAPG’24). T. Chakraborty acknowledges the support of the Rajiv Khemani Young Faculty Chair Professorship in Artificial Intelligence.
@@ Acknowledgements
While LORex effectively retrieves parent terms in taxonomy expansion, its performance heavily relies on the quality of the candidate ranking. If the correct parent term is not highly ranked, LORex suffers from increased retrieval errors and higher computational overhead due to multiple iterations of retrieval and verification. This iterative process significantly impacts both latency and accuracy, particularly in cases where the ranker fails to surface relevant candidates early. As a result, the overall system efficiency is constrained by the discriminative ranker, making LORex highly sensitive to ranking quality. In other words, “the method is only as good as its ranker.” Exploring better ranking techniques and reducing the reliance of retrievers on rankers will be the focus of our future works. We will also focus on improving interpretability by employing an ensemble of LLMs to reason over hierarchies. While this work focuses on taxonomy expansion, extending the framework to broader tasks such as taxonomy completion presents a promising direction for future research.
@@
Appendix A Modified Euler Tour
As discussed in Section 4.1, we present an example of modified Euler-toured path for taxonomy shown in Fig. 6. The Euler toured path is defined as $P$ = ["water pollution", "pollution", "environment", "pollution", "air pollution", "pollution", "soil pollution", "pollution", "water pollution", "marine pollution", "water pollution", "chemical pollution"]. Therefore, the two types of verbalized paths (with and without query definition are discussed in Prompt 1 as shown in Fig. 8, where $P\in P_{v}(q,P)$ while $P_{r}\in P_{v}(P)$. These paths are used to compute three types of losses as shown in Fig. 7.
@@ Appendix A Modified Euler Tour
Verbalized Paths $P$ potable water pollution is the contamination of water that is intended for human consumption [SEP] "water pollution" is child of "pollution" is child of "environment" is parent of "pollution" is parent of "air pollution" is child of "pollution" is parent of "soil pollution" is child of "pollution" is parent of "water pollution" is parent of "marine pollution" is child of "water pollution" is parent of "chemical pollution" $P_{r}$water pollution" is child of "pollution" is child of "environment" is parent of "pollution" is parent of "air pollution" is child of "pollution" is parent of "soil pollution" is child of "pollution" is parent of "water pollution" is parent of "marine pollution" is child of "water pollution" is parent of "chemical pollution"
Figure 8: Prompt 1 – Verbalized Path Examples of $P$ and $P_{r}$.
You are a semantic relevance expert for terms present in a taxonomy. Your task is to determine whether the set of candidate terms is a semantically relevant match for the given query term ’Arctic Ocean’ and ’Environment’ taxonomy.
List of Candidate terms:
- geophysical environment
- ocean
- wild mammal
- animal life
- climatic zone
Definitions of candidate terms:
- geophysical environment - environment most often refers to:
- ocean - an ocean is a body of water that composes much of a planet’s hydrosphere
- wild mammal - mammals, fur or hair, and three middle ear bones
- animal life - animals are multicellular eukaryotic organisms that form the biological kingdom animalia
- climatic zone - climate classification systems are ways of classifying the world’s climates
Reason over the definitions of candidate terms to determine their relevance as a whole with respect to query terms Arctic Ocean and Environment taxonomy. If at least one of the candidate terms is the most relevant parent term for the taxonomy, return YES. Otherwise, return NO.
Answer:
Figure 9: Prompt 2 – Semantic Filter Prompt Example
You are an expert in hypernymy (is-a) relationship detection for a taxonomy. Your task is to find the most appropriate candidate hypernym of the query node ’Arctic Ocean’ within the ’environment’ taxonomy. The most appropriate hypernym is the most granular category that directly encompasses the query term.
List of Candidate terms:
- geophysical environment
- ocean
- wild mammal
- animal life
- climatic zone
Definitions of candidate terms:
- geophysical environment - environment most often refers to:
- ocean - an ocean is a body of water that composes much of a planet’s hydrosphere
- wild mammal - mammals, fur or hair, and three middle ear bones
- animal life - animals are multicellular eukaryotic organisms that form the biological kingdom animalia
- climatic zone - climate classification systems are ways of classifying the world’s climates
Path from query term to root node for all candidate terms:
- Arctic Ocean -> geophysical environment -> environment
- Arctic Ocean -> ocean -> geophysical environment -> environment
- Arctic Ocean -> wild mammal -> animal life -> wildlife -> environment
- Arctic Ocean -> animal life -> wildlife -> environment
- Arctic Ocean -> climatic zone -> climate -> environment
Query Node Definition: The Arctic Ocean is the smallest and shallowest of the world’s oceans.
Some examples of hypernymy relationships in the taxonomy are as follows:
- Children of geophysical environment are: arid zone, estuary, island, lake, mountain, ocean, plain, polar region, sea
- Parent of animal life is: wildlife
- Children of climatic zone are: equatorial zone, frigid zone, humid zone, subtropical zone, temperate zone, tropical zone
- Parent of climatic zone is: climate
Instructions:
- Determine the most appropriate direct hypernym of ’Arctic Ocean’ from the given list of candidate terms. Do not return any term which is not in the list.
- A hypernym must be a most granular category in which the query node is an instance.
- If no candidate term correctly fits as a hypernym , return NOT FOUND.
- Do not include explanations, justifications, or additional context.
Answer:
Figure 10: Prompt 3 – Parent Retriever Prompt Example
You are an expert verifier of hypernymy relationship for a taxonomy using paths. You have been given the following path from query term ’Arctic Ocean’ to root node: Arctic Ocean -> ocean -> geophysical environment -> environment
Your task is to verify if the given path is the most appropriate path for the query node or the following paths provide a better alternative.
Other possible paths:
- Arctic Ocean -> geophysical environment -> environment
- Arctic Ocean -> ocean -> geophysical environment -> environment
- Arctic Ocean -> wild mammal -> animal life -> wildlife -> environment
- Arctic Ocean -> animal life -> wildlife -> environment
- Arctic Ocean -> climatic zone -> climate -> environment
Some examples of hypernymy relationships in the taxonomy are as follows:
- Children of geophysical environment are: arid zone, estuary, island, lake, mountain, ocean, plain, polar region, sea
- Parent of animal life is: wildlife
- Children of climatic zone are: equatorial zone, frigid zone, humid zone, subtropical zone, temperate zone, tropical zone
- Parent of climatic zone is: climate
Definitions of candidate terms:
- geophysical environment - environment most often refers to:
- ocean - an ocean is a body of water that composes much of a planet’s hydrosphere
- wild mammal - mammals, fur or hair, and three middle ear bones
- animal life - animals are multicellular eukaryotic organisms that form the biological kingdom animalia
- climatic zone - climate classification systems are ways of classifying the world’s climates
Query Node Definition: The Arctic Ocean is the smallest and shallowest of the world’s oceans.
Instructions: - Compare the given path with the alternative paths based on definitions and hypernymy relationships.
- Select the most appropriate path that best represents the hierarchical relationship of ’Arctic Ocean’ in the taxonomy.
Return the most appropriate path.
Answer:
Figure 11: Prompt 4 – Parent Verifier Prompt Example
@@ Appendix B Semantic Filter Prompt
The semantic filter module, discussed in Section 4.2, relies on LLM to check if the chunk of candidates is relevant by reasoning over their definitions. The prompt example is discussed in Prompt 2 shown in Fig. 9.
@@ Appendix C Parent Retrieval Prompt
The parent retriever module, discussed in Section 4.3, reasons over taxonomy paths to determine the appropriate parent from the chunk of candidate terms. An example of the parent retriever prompt is discussed in Prompt 3 as shown in Fig.  10.
@@ Appendix D Parent Verifier Prompt
The parent verifier module verifies the appropriateness of the retrieved candidate by reasoning on paths as discussed in Section 4.4. An example of the prompt is discussed in Prompt 4 as shown in Fig. 11.
Hyperparameter Sci Food Env WordNet
Batch Size 32 32 64 32
Learning Rate 2e-5 1e-5 2e-5 1e-5
Max Padding 350 350 200 350
$d$ 0.2 0.2 0.2 0.2
$\epsilon$ 1e-8 1e-8 1e-8 1e-8
Epochs 20 50 20 20
Table 7: Hyperparameter settings for different datasets.
@@ 14      $\mathcal{B}\leftarrow\mathcal{B}\setminus B^{i}_{j}$;
Appendix E Pseudocode
@@ Appendix F Implementation Details
LORex  and BERT+MLP are implemented in PyTorch, with all other baselines obtained from their respective official repositories. The training and inference processes are executed on a single 80GB NVIDIA A100 GPU to ensure computational efficiency. To implement TEMPORA, we use bert-base-uncased as the default pre-trained model, optimized using AdamW. The hyperparameters utilized for training and inference are detailed in Table 7. Notably, TEMPORA is trained on all benchmarks for fewer epochs compared to TEMP, as the primary objective is to achieve a ranker that delivers sufficiently high performance without excessive fine-tuning. Retrieval and verification rely on instruction-tuned models from the LLaMA family, namely LLaMA-3.1-8B-Instruct, LLaMA-3-8B-Instruct, and LLaMA-2-7B-Chat. As demonstrated in Fig. 12, all benchmarks except Food achieve a Hit@15 score above 0.9, indicating that for 90% of queries the correct parent exists within the top-15 candidates. Therefore, to enhance the efficiency of retrieval and verification, inference is performed on three chunks or 15 candidates simultaneously. If any chunk is empty, it is discarded, and only the remaining chunks are considered. This strategy ensures that the correct output is obtained in a single inference pass, optimizing both computational efficiency and retrieval accuracy. For text generation tasks, a stopping criterion is enforced to terminate generation upon encountering a newline character, which prevents the generation of unnecessary tokens, thereby reducing inference latency.
@@
Method Food WordNet
Hit@1 Hit@10 Hit@1 Hit@10
BM25 0.071 0.165 0.065 0.171
SBERTCOS 0.182 0.380 0.113 0.390
BERTCOS 0.077 0.118 0.076 0.225
SimCSE 0.119 0.280 0.258 0.621
TEMP 0.330 0.529 0.329 0.682
TEMPORA 0.451 0.663 0.405 0.794
Query True Parent Batch Skipped
fire protection environmental protection [environmental protection, environment, nuisance, environmental policy, pollution control measures]
desert geophysical environment [environment, degradation of the environment, geophysical environment, physical environment, wildlife]
demography sociology [social science, sociology, anthropology, economics, correlation]
thanatology science [endocrinology, anesthesiology, science, internal medicine, podiatry]
eugenics genetics [biology, genetics, life science, medical science, molecular biology]
scrumpy cider [cider, hard cider, sweet cider, alcohol, hooch]
Table 10: Failure cases where the LLM-based semantic filter incorrectly filters out batches with true parent.
@@ Appendix H Baseline Methods
  • FLAME Mishra et al. (2024) fine-tunes low-rank parameters of LLMs to predict parent terms by leveraging semantically relevant local and global neighborhood information of candidates.
Query Parent Retriever (Ground Truth) Path-Based Verifier Candidate-Based Verifier
@@
candied citrus peel (candied citrus peel)
@@
demography
the scientific study of human populations, including their size, composition, distribution, density, and growth.
sociology (sociology)
The most appropriate direct hypernym of ’demography’ is ’sociology.’
demography > sociology > social science > science
The path is appropriate because ’demography’ is a subfield of sociology, which belongs to social science, a domain within science. This path reflects a logical hierarchical classification of ’demography’.
social science
Although social science is a broader category that includes demography, the term ’sociology’ is more specific and directly connected to the study of population structures. Therefore, ’sociology’ better captures the essence of the query than ’social science’.
@@
differential calculus
a branch of calculus concerned with the study of the rates at which quantities change.
calculus (calculus)
The most appropriate direct hypernym of ’differential calculus’ is ’calculus’.
differential calculus > calculus > pure mathematics > mathematics > science
The path traverses through calculus and correctly situates the concept in the broader scientific hierarchy.
calculus
The candidate calculus is appropriate.
destruction of crops
a crop is a plant or animal product that can be grown and harvested for profit or subsistence. Destruction refers to damage to this resource.
degradation of environment (degradation of environment)
The most appropriate direct hypernym of ’destruction of crops’ is ’degradation’ of environment.
destruction of crops > degradation of the environment > environment
This path is appropriate because degradation of the environment includes anthropogenic harm to agricultural systems.
degradation of environment
The candidate node degradation of environment is valid.
Table 11: Case study analyzing the impact of LORex’s components on "Env", "Food" and "Science" benchmarks.
We adapt taxonomy completion baselines, such as TaxoComplete and TacoPrompt for the taxonomy expansion setting by introducing a pseudo-placeholder in place of the child node. This transformation ensures that the task is framed as a leaf node insertion, allowing these completion methods, originally designed for intermediate node prediction, to be effectively applied to taxonomy expansion tasks.
Appendix I Evaluation Metrics
    where $\mathbb{I}(\cdot)$ denotes the indicator function.
  • Wu & Palmer Similarity (Wu&P): This metric quantifies the structural similarity between the predicted parent and the ground truth within the seed taxonomy. It is computed based on the closest common ancestor,
[…]
Appendix J Discriminator Results
Appendix K Failure Cases of Semantic Filtering
While the semantic filter module is effective in improving efficiency and recall (as discussed in Section 5.3), there are a few failure cases where it incorrectly filters out batches containing the true parent term. These errors typically stem from ambiguity or insufficient granularity in the definitions used for semantic matching. As illustrated in Table 10, for the query term fire protection, the filter erroneously rejects the batch containing the correct parent environmental protection, with the LLM reasoning that fire protection does not semantically fall under that category. Similar misclassifications are observed for terms like desert and thanatology, where broader parent categories such as geophysical environment and science were excluded due to narrow or domain-specific interpretations. These cases highlight the limitations of LLM-based semantic inference when reasoning over abstract or generalized definitions, though such instances remain infrequent across datasets.
Appendix L Case Studies
Extending on the case studies discussed in Section 5.4, we further analyze some additional examples with more semantically rich explanations in Table 11. The parent retriever provides straightforward reasoning over definitions and path-based relationships. However, as it is susceptible to errors, we rely on the verifier’s reasoning to validate and refine the retriever’s predictions. The path-based verifier employs lineage-oriented reasoning to assess retrieval accuracy by analyzing hierarchical paths. For instance, in the case of ‘grapefruit peel’, the verifier explains, “The given path is the most appropriate as it directly connects ‘grapefruit peel’ to the root node ‘food’ through a series of hypernymy relationships that align with its definition.” In contrast, the candidate-based verifier relies on definitions, which can lead to incorrect predictions when contextual hierarchical relationships are not explicitly captured in definitions.